What is the combination of C (20,3)?
Answer by Maths68(1474)

=C(20,3)
=20!/3!(20-3)!
=20!/3!17!
=(20*19*18*17!)/(3*2*1*17!)
Cancel out 17!
=(20*19*18)/6
=20*19*3
=1140